# The Complexity of Sequential Prediction in Dynamical Systems

### v1

@@ Abstract
We study the problem of learning to predict the next state of a dynamical system when the underlying evolution function is unknown. Unlike previous work, we place no parametric assumptions on the dynamical system, and study the problem from a learning theory perspective. We define new combinatorial measures and dimensions and show that they quantify the optimal mistake and regret bounds in the realizable and agnostic setting respectively.
@@ 1 Introduction
Although we allow the state space $\mathcal{X}$ to be arbitrary, we only consider the 0-1 loss, which may be more appropriate for discrete-space dynamical systems. However, even discrete-space dynamical systems can be very expressive, capturing complex natural processes (Hoekstra et al., 2010). One important example of such dynamical systems is cellular automata (Wolfram, 1986). A $d$-dimensional boolean cellular automata is defined on a state space $\mathcal{X}=\{0,1\}^{\mathbb{Z}^{d}}$, the set of $\{0,1\}$-valued $d$-dimensional square grids defined using $d$-dimensional integer lattices. Square grids are called the cells and the boolean values in the cells are referred to as automaton. The evolution function $f$ is defined using a local updating rule $h:\{0,1\}\times\{0,1\}^{n}\to\{0,1\}$, where the automaton of the cell at time $t+1$ depends on the automaton of the cell as well as the automata of all the neighboring cells at time $t$. Here, $n$ is the number of neighboring cells. Since $n=3^{d}-1$, there are $2^{3^{d}}$ possible local updating rules and thus evolution functions for $d$-dimensional boolean cellular automata. Since the pioneering work of (von Neumann, 1966), who used it to model biological self-reproduction, cellular automata have been widely used in complex systems modeling (Batty, 2005; Barberousse and Imbert, 2013; Xiao et al., 2011).
Given any learning problem $(\mathcal{X},\mathcal{F})$, we aim to find necessary and sufficient conditions for the learnability of $\mathcal{F}$ while also quantifying the minimax rates under two notions of expected regret: Markovian and Flow regret. To that end, our main contributions are summarized below.
@@
    We provide a qualitative and sharp quantitative characterization of learnability in the realizable setting, when there exists an evolution function in the class $\mathcal{F}$ that generates the sequence of states. Our characterization is in terms of two new combinatorial objects: the Evolution complexity and Evolution dimension. We use these complexity measures to compute the minimax expected mistakes for three classes of linear dynamical systems.
@@
    We show that all rates are possible for the minimax expected mistakes when learning dynamical systems in the realizable setting. This is in contrast to online multiclass classification where only two rates are possible. Nevertheless, we show that a combinatorial parameter termed the Branching dimension characterizes constant minimax expected mistakes.
  • (iii)
    In the agnostic setting under Markovian regret, we show that the Littlestone dimension provides a characterization of learnability. As a corollary, we establish a separation between realizable and agnostic learnability under Markovian regret.
  • (iv)
    We show that the separation between realizable and agnostic learnability continues to hold when considering Flow regret. However, if the evolution class has uniformly bounded projections, then we show that realizable and agnostic learnability under Flow regret are equivalent.
For (i), our bounds on the minimax expected mistakes in terms of the Evolution complexity are constructive. For the upperbound, we construct a minimax optimal algorithm. For the lowerbound, we construct a hard trajectory. For (ii), given any rate, we construct an evolution class whose minimax expected mistakes matches this rate. To prove (iii), our bounds on the minimax expected regret are also constructive. For the upperbound, we show that the Littlestone dimension is sufficient by reducing learning dynamical systems to online multiclass classification. To show that the Littlestone dimension is necessary, we also construct a hard trajectory. To establish the separation between realizable and agnostic learnability, we give a class that is learnable in the realizable setting but not in the agnostic setting under Markovian regret.
@@ 1.1 Related Works
There has been a long line of work studying prediction and regret minimization when learning unknown dynamical systems (Hazan et al., 2017, 2018; Ghai et al., 2020; Lee, 2022; Rashidinejad et al., 2020; Kozdoba et al., 2019; Tsiamis and Pappas, 2022; Lale et al., 2020). However, these works focus on prediction for fully/partially-observed linear dynamical systems under various data-generating processes. Moreover, there is also a line of work on regret minimization for linear dynamical systems for online control problems (Abbasi-Yadkori and Szepesvári, 2011; Cohen et al., 2018; Agarwal et al., 2019). For non-linear dynamical systems, there has been some applied work studying data-driven approaches to prediction (Wang et al., 2016; Korda and Mezić, 2018; Ghadami and Epureanu, 2022). Regret minimization for non-linear dynamical systems has mainly been studied in the context of control (Kakade et al., 2020; Muthirayan and Khargonekar, 2021).
Another important line of work is that of system identification and parameter estimation (Åström and Eykhoff, 1971; Ljung, 1999). Here, the goal is to recover and estimate the parameters of the underlying evolution function from the observed sequence of states. There is a long history of work studying system identification in both the batch (Campi and Weyer, 2002; Vidyasagar and Karandikar, 2006; Foster et al., 2020; Sattar and Oymak, 2022; Bahmani and Romberg, 2020) and streaming settings (Kowshik et al., 2021a, b; Giannakis et al., 2023; Jain et al., 2021). Several works have also considered the problem of learning the unknown evolution rule of dynamical systems defined over discrete state spaces. For example, Wulff and Hertz (1992) train a neural network on a sequence of observed states to approximate the unknown evolution rule of a cellular automaton. Grattarola et al. (2021) extend this work to learning the unknown evolution rule of a graph cellular automaton, a generalization of a regular cellular automaton, using graph neural networks. (Qiu et al., ; Rosenkrantz et al., 2022) also consider the problem of PAC learning a discrete-time, finite-space dynamical system defined over a (un)directed graph.
A related direction to our work is online learning for time series forecasting, where autoregressive models are used to predict the next state (Anava et al., 2013, 2015; Liu et al., 2016; Yang et al., 2018). Finally, there is a rich history of characterizing learnability in terms of complexity measures and combinatorial dimensions in learning theory (Vapnik and Chervonenkis, 1971; Littlestone, 1987; Natarajan, 1989; Kearns and Schapire, 1994; Bartlett and Mendelson, 2002; Daniely et al., 2011; Daniely and Shalev-Shwartz, 2014; Rakhlin et al., 2015).
@@ 2.1 Discrete-time Dynamical Systems
A discrete-time dynamical system is a tuple $(\mathbb{N},\mathcal{X},f)$, where $\mathbb{N}$ is the set of natural numbers denoting the time steps, $\mathcal{X}$ is a non-empty set called the state space, and $f:\mathcal{X}\to\mathcal{X}$ is a deterministic map that defines the evolution of the dynamical system. That is, the $(t+1)$-th iterate of the dynamics can be expressed in terms of $t$-th iterate using the relation $x_{t+1}=f(x_{t})$. Define $f^{t}$ to be the $t$-fold composition of $f$. That is, $f^{2}=f\circ f$, $f^{3}=f\circ f\circ f$, and so forth. Given an initial state $x_{0}\in\mathcal{X}$, the sequence $\{f^{t}(x_{0})\}_{t\in\mathbb{N}}$ is called the flow of the dynamical system through $x_{0}$. Finally, let $\mathcal{F}\subseteq\mathcal{X}^{\mathcal{X}}$ denote a class of evolution functions on the state space $\mathcal{X}$ and $\mathcal{F}(x)=\{f(x)\,\mid f\in\mathcal{F}\}\subseteq\mathcal{X}$ to be the projection of $\mathcal{F}$ onto $x\in\mathcal{X}$.
@@ 2.2 Learning-to-Predict in Dynamical Systems
$\texttt{R}^{\texttt{MK}}_{\mathcal{A}}(T,\mathcal{F}):=\sup_{(x_{0},x_{1},\ldots,x_{T})}\mathbb{E}\left[\sum_{t=1}^{T}\mathbbm{1}\{\mathcal{A}(x_{<t})\neq x_{t}\}\right]-\inf_{f\in\mathcal{F}}\sum_{t=1}^{T}\mathbbm{1}\{f(x_{t-1})\neq x_{t}\},$ (1)
@@
Definition 1 (Agnostic Learnability under Markovian Regret).
An evolution class $\mathcal{F}\subseteq\mathcal{X}^{\mathcal{X}}$ is learnable in the agnostic setting if and only if $\,\,\,\inf_{\mathcal{A}}\emph{{R}}^{\emph{{MK}}`}_{\mathcal{A}}(T,\mathcal{F})=o(T).$
@@
$\texttt{M}_{{\mathcal{A}}}(T,\mathcal{F}):=\sup_{x_{0}}\,\sup_{f\in\mathcal{F}}\,\mathbb{E}\left[\sum_{t=1}^{T}\mathbbm{1}\{\mathcal{A}(x_{<t})\neq f(x_{t-1})\}\right].$
@@
Definition 2 (Realizable Learnability).
An evolution class $\mathcal{F}\subseteq\mathcal{X}^{\mathcal{X}}$ is learnable in the realizable setting if and only if $\,\,\,\inf_{\mathcal{A}}\emph{{M}}_{{\mathcal{A}}}(T,\mathcal{F})=o(T).$
@@ 2.3 Complexity Measures
Figure 1: Shattered Trajectory (left) and Littlestone (right) trees of depth $2$.
@@
Definition 3 (Trajectory tree).
@@
See Figure 1 for an example of a trajectory tree. Given a trajectory tree $\mathcal{T}$ of depth $d$, a root-to-leaf path down $\mathcal{T}$ is defined by a string $\sigma\in\{-1,1\}^{d}$ indicating whether to go left ($\sigma_{t}=-1$) or to go right ($\sigma_{i}=+1$) at each depth $t\in[d]$. A path $\sigma\in\{-1,1\}^{d}$ down $\mathcal{T}$ gives a trajectory $\{x_{t}\}_{t=0}^{d}$, where $x_{0}$ denotes the instance labeling the root node and $x_{t}$ is the instance labeling the edge following the prefix $(\sigma_{1},...,\sigma_{t})$ down the tree. A path $\sigma\in\{-1,1\}^{d}$ down $\mathcal{T}$ is shattered by $\mathcal{F}$ if there exists a $f\in\mathcal{F}$ such that $f(x_{t-1})=x_{t}$ for all $t\in[d]$, where $\{x_{t}\}_{t=0}^{d}$ is the corresponding trajectory obtained by traversing $\mathcal{T}$ according to $\sigma$. If every path down $\mathcal{T}$ is shattered by $\mathcal{F}$, we say that $\mathcal{T}$ is shattered by $\mathcal{F}$.
@@
Definition 4 (Branching factor).
@@
$\emph{{B}}(\mathcal{T}):=\min_{\sigma\in\{-1,1\}^{d}}\,\sum_{t=1}^{d}\mathbbm{1}\{\mathcal{T}_{t}((\sigma_{<t},-1))\neq\mathcal{T}_{t}\big{(}(\sigma_{<t},+1)\big{)}\}.$
The branching factor of a path $\sigma\in\{-1,1\}^{d}$ captures the distinctness of states labeling the two children of internal nodes in this path. In particular, it counts the number of nodes in the path whose two children are labeled by distinct states. The branching factor of a trajectory tree is just the smallest branching factor across all paths. Using the notion of shattering and Definition 4, we define a new complexity measure, termed the Evolution complexity, of a function class $\mathcal{F}$.
Definition 5 (Evolution complexity).
Let $\mathcal{S}(\mathcal{F},d)$ be the set of all trajectory trees of depth $d\in\mathbb{N}$ shattered by $\mathcal{F}$. Then, the Evolution complexity of $\mathcal{F}$ at depth $d$ is defined as $\emph{{C}}_{d}(\mathcal{F}):=\sup_{\mathcal{T}\in\mathcal{S}(\mathcal{F},d)}\emph{{B}}(\mathcal{T}).$
In Section 3, we show that the Evolution complexity exactly (up to a factor of $2$) captures the minimax expected mistakes in the realizable setting. We note that there is an existing notion of complexity for dynamical systems, termed topological entropy, that quantifies the complexity of a particular evolution function $f\in\mathcal{F}$ (Adler et al., 1965). However, topological entropy does not characterize learnability as =$\mathcal{F}=\{f\}$ is trivially learnable when $f$ has infinite topological entropy.

### v2

@@ Abstract
We study the problem of learning to predict the next state of a dynamical system when the underlying evolution function is unknown. Unlike previous work, we place no parametric assumptions on the dynamical system, and study the problem from a learning theory perspective. We define new combinatorial measures and dimensions and show that they quantify the optimal mistake and regret bounds in the realizable and agnostic settings respectively. By doing so, we find that in the realizable setting, the total number of mistakes can grow according to any increasing function of the time horizon $T$. In contrast, we show that in the agnostic setting under the commonly studied notion of Markovian regret, the only possible rates are $\Theta(T)$ and $\tilde{\Theta}(\sqrt{T})$.
keywords:
Discrete-time Dynamical Systems, Online Learnability
@@ 1 Introduction
Although we allow the state space $\mathcal{X}$ to be arbitrary, we only consider the 0-1 loss, which may be more appropriate for discrete-space dynamical systems. However, even discrete-space dynamical systems can be very expressive, capturing complex processes like cellular automata (Hoekstra et al., 2010; Wolfram, 1986) and language modeling (Elman, 1995). For example, let $\mathcal{V}$ be a countable token space and $\mathcal{X}=\mathcal{V}^{\star}$ to be the state space containing all finite sequences of elements from $\mathcal{V}$. Consider a function class $\mathcal{F}\subseteq\mathcal{X}^{\mathcal{X}}$ with the following property: for every $f\in\mathcal{F}$ and any input $x\in\mathcal{X}$, $f(x)=x\circ v$, where we use $\circ$ to represent the concatenation operator. Then, $\mathcal{F}$ is a set of auto-regressive models which given a sequence of tokens $x\in\mathcal{X}$, predicts the next token $v\in\mathcal{V}$ in the sequence. In particular, $\mathcal{F}$ could be a class of language models which use deterministic decoding strategies (e.g. greedy decoding) to output the next token (Chorowski and Jaitly, 2016). For such a function class $\mathcal{F}$, one might be interested in understanding the optimal number of mistakes for next-token prediction when the true sequence of tokens is generated by some unknown $f\in\mathcal{F}$.
Given any learning problem $(\mathcal{X},\mathcal{F})$, we aim to find necessary and sufficient conditions for the learnability of $\mathcal{F}$ while also quantifying the minimax rates under two notions of expected regret: Markovian and Flow regret (Equations 1 and 2 respectively). To that end, our main contributions are summarized below.
@@
    We provide a quantitative characterization of learnability in the realizable setting, when there exists an evolution function in the class $\mathcal{F}$ that generates the sequence of states. Our characterization is in terms of a new combinatorial complexity measure we call the Evolution complexity. Using this characterization, we show that all rates are possible for the minimax expected mistakes when learning dynamical systems in the realizable setting. This is in contrast to online multiclass classification where only two rates are possible. Finally, we compare realizable learnability of dynamical systems to realizable learnability in PAC and online classification.
@@
    In the agnostic setting, we lower and upper bound the minimax Markovian regret in terms of the Littlestone dimension. This result shows that the finiteness of the Littlestone dimension characterizes learnability under Markovian regret. As a corollary, we establish a separation between realizable and agnostic learnability under Markovian regret. We show this separation between realizable and agnostic learnability continues to hold when considering Flow regret. However, if the evolution class has uniformly bounded projections, we show that realizable and agnostic learnability under Flow regret are equivalent.
Our characterization of realizable learnability in terms of the Evolution complexity follows from standard techniques in online classification. However, our results showing all possible rates requires a careful construction of a family of classes which have not been studied in learning theory. Likewise, our comparisons of realizable learnability require a careful construction of non-trivial classes, and computing their combinatorial dimensions.
In the agnostic setting, our upper bound on the minimax Markovian regret in terms of the Littlestone dimension results from reducing learning dynamical systems to online multiclass classification. However, both of the lower bounds $\Omega(\sqrt{T})$ and $\Omega(\operatorname{L}(\mathcal{H}))$ in Theorem 4.1 are not standard. The lower bound of $\Omega(\operatorname{L}(\mathcal{H}))$ requires constructing a hard stream by traversing down a Littlestone tree skipping certain levels. The lower bound of $\sqrt{T}$ requires constructing a hard stream using two different evolution functions $f_{1},f_{2}$ by: (a) setting $x_{0}$ to be a state they differ on (b) generating their trajectories starting from $x_{0}$ and finally (c) picking states from each trajectory in an alternating fashion. These arguments are different from the typical lower bound construction for online classification. In addition, the construction in Theorem 4.4 showing the separation between realizable and agnostic learnability under Flow regret is non-trivial. It involves constructing an $\mathcal{F}\subseteq\mathcal{X}^{\mathcal{X}}$ so that on a large subset of states in $\mathcal{X}$, every function $f\in\mathcal{F}$ effectively reveals itself.
@@ 1.1 Related Works
There has been a long line of work studying prediction and regret minimization when learning unknown dynamical systems (Hazan et al., 2017, 2018; Ghai et al., 2020; Lee, 2022; Rashidinejad et al., 2020; Kozdoba et al., 2019; Tsiamis and Pappas, 2022; Lale et al., 2020). However, these works focus on prediction for fully/partially-observed linear dynamical systems under various data-generating processes. Moreover, there is also a line of work on regret minimization for linear dynamical systems for online control problems (Abbasi-Yadkori and Szepesvári, 2011; Cohen et al., 2018; Agarwal et al., 2019). For non-linear dynamical systems, there has been some applied work studying data-driven approaches to prediction (Wang et al., 2016; Korda and Mezić, 2018; Ghadami and Epureanu, 2022). Regret minimization for non-linear dynamical systems has mainly been studied in the context of control (Kakade et al., 2020; Muthirayan and Khargonekar, 2021). Another related line of work is online learning for time series forecasting, where autoregressive models are used to predict the next state (Anava et al., 2013, 2015; Liu et al., 2016; Yang et al., 2018).
Another important line of work is that of system identification and parameter estimation (Åström and Eykhoff, 1971; Ljung, 1999). Here, the goal is to recover and estimate the parameters of the underlying evolution function from the observed sequence of states. There is a long history of work studying system identification in both the batch (Campi and Weyer, 2002; Vidyasagar and Karandikar, 2006; Foster et al., 2020; Sattar and Oymak, 2022; Bahmani and Romberg, 2020) and streaming settings (Kowshik et al., 2021a, b; Giannakis et al., 2023; Jain et al., 2021). Several works have also considered the problem of learning the unknown evolution rule of dynamical systems defined over discrete state spaces. For example, Wulff and Hertz (1992) trains a neural network on a sequence of observed states to approximate the unknown evolution rule of a cellular automaton. Grattarola et al. (2021) extends this work to learning the unknown evolution rule of a graph cellular automaton, a generalization of a regular cellular automaton, using graph neural networks. (Qiu et al., ; Rosenkrantz et al., 2022) also consider the problem of PAC learning a discrete-time, finite-space dynamical system defined over a (un)directed graph. We also note the related work of Berry and Das (2023, 2025), who study the problem of learning dynamics observed through a continuous embedding and derive learning guarantees based on various structural properties of the underlying system. Finally, our work builds on a rich tradition in learning theory that characterizes learnability through complexity measures and combinatorial dimensions (Vapnik and Chervonenkis, 1971; Littlestone, 1987; Bartlett and Mendelson, 2002; Daniely et al., 2011).
@@ 2.1 Discrete-time Dynamical Systems
A discrete-time dynamical system is a tuple $(\mathbb{N},\mathcal{X},f)$, where $\mathbb{N}$ is the set of natural numbers denoting the time steps, $\mathcal{X}$ is a non-empty set called the state space. In this work, we make no assumption on the cardinality of $\mathcal{X}$, so it can be unbounded and perhaps even uncountable. The function $f:\mathcal{X}\to\mathcal{X}$ is a deterministic map that defines the evolution of the dynamical system. That is, the $(t+1)$-th iterate of the dynamics can be expressed in terms of $t$-th iterate using the relation $x_{t+1}=f(x_{t})$. Define $f^{t}$ to be the $t$-fold composition of $f$. That is, $f^{2}=f\circ f$, $f^{3}=f\circ f\circ f$, and so forth. Given an initial state $x_{0}\in\mathcal{X}$, the sequence $\{f^{t}(x_{0})\}_{t\in\mathbb{N}}$ is called the flow of the dynamical system through $x_{0}$. Finally, let $\mathcal{F}\subseteq\mathcal{X}^{\mathcal{X}}$ denote a class of evolution functions on the state space $\mathcal{X}$ and $\mathcal{F}(x)=\{f(x)\,\mid f\in\mathcal{F}\}\subseteq\mathcal{X}$ to be the projection of $\mathcal{F}$ onto $x\in\mathcal{X}$.
@@ 2.2 Learning-to-Predict in Dynamical Systems
$\operatorname{MR}_{\mathcal{A}}(T,\mathcal{F}):=\sup_{(x_{0},x_{1},\ldots,x_{T})}\mathbb{E}\left[\sum_{t=1}^{T}\mathbbm{1}\{\mathcal{A}(x_{<t})\neq x_{t}\}\right]-\inf_{f\in\mathcal{F}}\sum_{t=1}^{T}\mathbbm{1}\{f(x_{t-1})\neq x_{t}\},$ (1)
@@
Definition 2.1 (Agnostic Learnability under Markovian Regret).
[…]
Perhaps a more natural definition of expected regret in the agnostic setting is to compare the prediction of the learner to the prediction of the best-fixed trajectory generated by functions in our evolution class. To that end, define
$\operatorname{FR}_{\mathcal{A}}(T,\mathcal{F}):=\sup_{(x_{0},x_{1},\ldots,x_{T})}\Bigl{(}\mathbb{E}\left[\sum_{t=1}^{T}\mathbbm{1}\{\mathcal{A}(x_{<t})\neq x_{t}\}\right]-\inf_{f\in\mathcal{F}}\sum_{t=1}^{T}\mathbbm{1}\{f^{t}(x_{0})\neq x_{t}\}\Bigl{)}.$ (2)
Definition 2.2 (Agnostic Learnability under Flow Regret).
An evolution class $\mathcal{F}\subseteq\mathcal{X}^{\mathcal{X}}$ is learnable in the agnostic setting under Flow regret if and only if $\,\,\inf_{\mathcal{A}}\,\operatorname{FR}_{\mathcal{A}}(T,\mathcal{F})=\text{o}(T).$
@@
$\operatorname{M}_{{\mathcal{A}}}(T,\mathcal{F}):=\sup_{x_{0}}\,\sup_{f\in\mathcal{F}}\,\mathbb{E}\left[\sum_{t=1}^{T}\mathbbm{1}\{\mathcal{A}(x_{<t})\neq f(x_{t-1})\}\right].$
@@
Definition 2.3 (Realizable Learnability).
An evolution class $\mathcal{F}\subseteq\mathcal{X}^{\mathcal{X}}$ is learnable in the realizable setting if and only if $\,\,\,\inf_{\mathcal{A}}\operatorname{M}_{{\mathcal{A}}}(T,\mathcal{F})=o(T).$
@@ 2.3 Complexity Measures
Definition 2.4 (Trajectory tree).
@@
Given a trajectory tree $\mathcal{T}$ of depth $d$, a root-to-leaf path down $\mathcal{T}$ is defined by a string $\sigma\in\{-1,1\}^{d}$ indicating whether to go left ($\sigma_{t}=-1$) or to go right ($\sigma_{i}=+1$) at each depth $t\in[d]$. A path $\sigma\in\{-1,1\}^{d}$ down $\mathcal{T}$ gives a trajectory $\{x_{t}\}_{t=0}^{d}$, where $x_{0}$ denotes the instance labeling the root node and $x_{t}$ is the instance labeling the edge following the prefix $(\sigma_{1},...,\sigma_{t})$ down the tree. A path $\sigma\in\{-1,1\}^{d}$ down $\mathcal{T}$ is shattered by $\mathcal{F}$ if there exists a $f\in\mathcal{F}$ such that $f(x_{t-1})=x_{t}$ for all $t\in[d]$, where $\{x_{t}\}_{t=0}^{d}$ is the corresponding trajectory obtained by traversing $\mathcal{T}$ according to $\sigma$. If every path down $\mathcal{T}$ is shattered by $\mathcal{F}$, we say that $\mathcal{T}$ is shattered by $\mathcal{F}$.
@@
Definition 2.5 (Branching factor).
@@
$\operatorname{B}(\mathcal{T}):=\min_{\sigma\in\{-1,1\}^{d}}\,\sum_{t=1}^{d}\mathbbm{1}\{\mathcal{T}_{t}((\sigma_{<t},-1))\neq\mathcal{T}_{t}\big{(}(\sigma_{<t},+1)\big{)}\}.$
The branching factor of a path $\sigma\in\{-1,1\}^{d}$ captures the distinctness of states labeling the two children of internal nodes in this path. In particular, it counts the number of nodes in the path whose two children are labeled by distinct states. The branching factor of a trajectory tree is just the smallest branching factor across all paths. Using the notion of shattering and Definition 2.5, we define a new complexity measure, termed the Evolution complexity, of a function class $\mathcal{F}$.
Definition 2.6 (Evolution complexity).
Let $\mathcal{S}(\mathcal{F},d)$ be the set of all trajectory trees of depth $d\in\mathbb{N}$ shattered by $\mathcal{F}$. Then, the Evolution complexity of $\mathcal{F}$ at depth $d$ is defined as $\operatorname{C}_{d}(\mathcal{F}):=\sup_{\mathcal{T}\in\mathcal{S}(\mathcal{F},d)}\operatorname{B}(\mathcal{T}).$
In Section 3, we show that the Evolution complexity exactly (up to a factor of $2$) captures the minimax expected mistakes in the realizable setting. We provide some examples of classes $\mathcal{F}$ and their evolution complexities in Theorem 3.6. We note that there is an existing notion of complexity for dynamical systems, termed topological entropy, that quantifies the complexity of a particular evolution function $f\in\mathcal{F}$ (Adler et al., 1965). However, topological entropy does not characterize learnability as =$\mathcal{F}=\{f\}$ is trivially learnable when $f$ has infinite topological entropy.